$12,305 is invested, part at 10% and the rest at 8%. If the interest earned from the amount invested at 10% exceeds the interest earned from the amount invested at 8% by 1025.30, how much is invested at each rate?
Answer by mananth(16946)

Part I 10.00% per annum ------------- x
Part II 8.00% per annum ------------ y
total investment 12305
Interest----- from greater part = 10.00% x
Interest from smaller part= 8.00% y
greater by 1025.3
10.00% x = 8.00% y + 1025.3
multiply by 100
10 x - -8 y = 102530 --------1
x+ 1 y = 12305 -------- 2
eliminate y
Multiply 1 by 1
Multiply 2 by 8
we get
10 x - -8 y = 102530
8 x - 8 y = 98440
18 x = 200970
/ 18
x= 11165
Plug the value of x in in 1
111650 -8 y = 102530
-8 y = 102530 -111650
-8 y = -9120
/ -8
y= 1140.00
$ 11165 at 10.00%
$ 1140.00 at 8.00%

CHECK
1116.5 = 91.20 + 1025.3
1116.50